On January 2, 2019, Twilight Hospital purchased a $106,000 special radiology scanner from Pharoah Inc. The scanner had a useful life of 4 years and was estimated to have no disposal value at the end of its useful life. The straight-line method of depreciation is used on this scanner. Annual operating costs with this scanner are $106,000.

Approximately one year later, the hospital is approached by Dyno Technology salesperson, Jacob Cullen, who indicated that purchasing the scanner in 2019 from Pharoah Inc. was a mistake. He points out that Dyno has a scanner that will save Twilight Hospital $26,000 a year in operating expenses over its 3-year useful life. Jacob notes that the new scanner will cost $110,000 and has the same capabilities as the scanner purchased last year. The hospital agrees that both scanners are of equal quality. The new scanner will have no disposal value. Jacob agrees to buy the old scanner from Twilight Hospital for $54,500.

If Twilight Hospital sells its old scanner on January 2, 2020, compute the gain or loss on the sale.
$_______________
Answer #1

Answer: Loss $25,000

Explanation:

Gain/Loss on sale=Sales price of asset-Book value of Old asset

Sales price=54,500

Book Value of old asset= Original Cost-accumulated depreciation

accumulated depreciation(straight line method)=(original cost-salvage)*used years/useful life=(106,000-0)1/4=26,500

Book value=106,000-26,500=79,500

Gain/loss=54500-79500=-25,000 (Loss)

If helpful, Thumbs UP please:)